Let us consider a quadratic polynomial f(x) such that equation f(x) = 5x - 15 has exactly one root, and equation f(x) = 6x - 18 has exactly one root. Find the maximum value of a discriminant of f(x).
1
Expert's answer
2020-11-29T19:25:30-0500

Let "f(x)=ax^2+bx+c" be a quadratic polynomial.

The equation "f(x)=5x-15" has exactly one root


"ax^2+bx+c=5x-15"

"ax^2+(b-5)x+(c+15)=0"

"D_1=(b-5)^2-4a(c+15)=0"

The equation "f(x)=6x-18" has exactly one root


"ax^2+bx+c=6x-18"

"ax^2+(b-6)x+(c+18)=0"

"D_2=(b-6)^2-4a(c+18)=0"

"(b-5)^2-(b-6)^2-4a(c+15-c-18)=0"

"a=-\dfrac{1}{12}(2b-11)"

"b^2-10b+25-4ac+5(2b-11)=0"

"4ac=b^2-30"


"D_0=b^2-4ac=b^2-(b^2-30)=30"

The maximum value of a discriminant of f(x) is 30.